OpenStudy (toxicsugar22):

Find the first three terms of the product in acending powers. (1-2x^2-x^4)(4+x^2-2x^4)

OpenStudy (anonymous):

well, why don't you multiply the two polynomials and post what you got?

OpenStudy (toxicsugar22):

ok

OpenStudy (toxicsugar22):

i got 4-x^2-2x^4-8x^4x^6-4x^4-x^6-2x^8 is that right and if it isnt caould you please correct me

OpenStudy (anonymous):

okay so you got that. why don't you add up the like terms and post it here. for example, -2x^4 - 4x^4 = -6x^4

OpenStudy (anonymous):

there should be a plus of minus sign between 8x^4 and x^6 in the polynomial you posted

OpenStudy (toxicsugar22):

so my answer is not right

OpenStudy (toxicsugar22):

ohh i see there is a minus between 8x^4 and x^6

OpenStudy (toxicsugar22):

4-x^2-2x^4-8x^4-x^6-4x^4-x^6-2x^8

OpenStudy (anonymous):

okay, but you have to add up the like terms. you see, a^2 +3 a^2 = 4a^2. similarly -2x^4 - 8x^4 = -10x^4

OpenStudy (anonymous):

and so on. so group all the terms together and write down your expression

OpenStudy (toxicsugar22):

i dont know but i got 4-x^2-14x^4-2x^6-2x^8

OpenStudy (toxicsugar22):

is that right

OpenStudy (anonymous):

yes, that seems to be right. all you have to do now is see what the first three terms are in ascending order of powers.

OpenStudy (toxicsugar22):

is it 4-x^2-14x^4

OpenStudy (anonymous):

yes.

OpenStudy (toxicsugar22):

ok thanks you

OpenStudy (toxicsugar22):

cana you help me with on more promblem

OpenStudy (anonymous):

although you should say, the first three terms are 4, -x^2 and -14x^4

OpenStudy (anonymous):

you are welcome. post your other problem and we will see how to solve it.

OpenStudy (toxicsugar22):

A large wooden box is 3 feet longer than it is wide, and its height is 2 feet shorter than its width. a. If the width of the box is w, write expressions for its lenghth and its height. b. write a plynomial for the volume of the box. c. Write a polynomial for the surface area of the box

OpenStudy (toxicsugar22):

do u get it or no

OpenStudy (toxicsugar22):

ok srry not that one

OpenStudy (anonymous):

yes, i got it. do you want the answer or not?

OpenStudy (toxicsugar22):

yes

OpenStudy (toxicsugar22):

thanks

OpenStudy (anonymous):

okay it says the box is 3 feet longer than its width. so length = ?

OpenStudy (toxicsugar22):

i dont know can u show me the firt time please

OpenStudy (toxicsugar22):

can u help me please

OpenStudy (anonymous):

if width is w, length = w + 3feet.

OpenStudy (anonymous):

so if height is 2 feet less than width, height = ?

OpenStudy (toxicsugar22):

2-w

OpenStudy (toxicsugar22):

or w-2

OpenStudy (anonymous):

yes, w -2 is correct. so width is w, length is w+3 and height is w-2

OpenStudy (anonymous):

what is the volume and surface area then?

OpenStudy (toxicsugar22):

im not sure surface area is( w-2)(w+3)

OpenStudy (toxicsugar22):

can u help me

OpenStudy (anonymous):

what is the volume of a cuboid?

OpenStudy (anonymous):

just tell me the general formula.

OpenStudy (toxicsugar22):

ok the surface formula

OpenStudy (toxicsugar22):

s=2lh+2wh+2lw

OpenStudy (toxicsugar22):

is that it

OpenStudy (anonymous):

yes. so express the surface area in terms of the height width and length of your cube.

OpenStudy (toxicsugar22):

so do i plug in the 3 and 2 for the height and the width

OpenStudy (anonymous):

your height is w -2 and length is w+3 and width is w. so plug those values.

OpenStudy (toxicsugar22):

so s=2(w+3)(w-2)+2(w)(w-2)+2(w+3)(w)

OpenStudy (toxicsugar22):

is it right

OpenStudy (anonymous):

yes

OpenStudy (toxicsugar22):

now do i distrubuite

OpenStudy (anonymous):

yes, you can do that

OpenStudy (toxicsugar22):

how do i muliply 2(w+3)(w-2)

OpenStudy (toxicsugar22):

can you just show me

OpenStudy (anonymous):

http://www.webmath.com/polymult.html

OpenStudy (toxicsugar22):

ok i got 6w^2+4w-12 is that right

OpenStudy (toxicsugar22):

is it right

OpenStudy (anonymous):

no

OpenStudy (toxicsugar22):

than can u help me

OpenStudy (toxicsugar22):

please

OpenStudy (anonymous):

i gave you the link, did you use it?

OpenStudy (toxicsugar22):

yes ill do it again

OpenStudy (anonymous):

okay.

OpenStudy (toxicsugar22):

igot the same thing

OpenStudy (toxicsugar22):

unless im doing something wrong

OpenStudy (toxicsugar22):

r u there

OpenStudy (anonymous):

yes, \[2(w+3)(w-2)=2(w \times w +3 \times w -2 \times w -2 \times 3)\]

OpenStudy (anonymous):

\[= 2(w ^{2}+w-6)=2w ^{2}+2w-12\]

OpenStudy (toxicsugar22):

is that it

OpenStudy (toxicsugar22):

is that the anwer to that question

OpenStudy (anonymous):

no, that is the multiplication of 2(w+3)(w-2) the surface area is 2 lw+2lh+2hw and the volume is lwh

OpenStudy (toxicsugar22):

ok can u help me step by step please

OpenStudy (toxicsugar22):

im so lost

OpenStudy (toxicsugar22):

can u help

OpenStudy (anonymous):

what do you need help with? I already explained everything to you.

OpenStudy (toxicsugar22):

no answering the suface area

OpenStudy (toxicsugar22):

please like how do u solve 2(w+3)(w-2)+2(w)(w-2)+2(w+3)(w)

OpenStudy (anonymous):

2w(w-2) = 2w^2-4w 2w(w+3) = 2w^2+6w. 2(w+3)(w-2) = 2w^2 +2w-12 add all three and post what you get.

OpenStudy (toxicsugar22):

6w^4w-12 is that right

OpenStudy (toxicsugar22):

is it

OpenStudy (toxicsugar22):

please help me

OpenStudy (toxicsugar22):

r u there

OpenStudy (anonymous):

yes that is right. that is the surface area.

OpenStudy (toxicsugar22):

ok and the volume of the box

OpenStudy (toxicsugar22):

the surface area of the the box is 6w^6+4w-12

OpenStudy (anonymous):

how did you get 6w^6 term in there?

OpenStudy (toxicsugar22):

i dk

OpenStudy (anonymous):

can you add the three terms like I said and post it here properly?

OpenStudy (toxicsugar22):

i did

OpenStudy (anonymous):

2w(w-2) = 2w^2-4w 2w(w+3) = 2w^2+6w. 2(w+3)(w-2) = 2w^2 +2w-12 add all three and post what you get

OpenStudy (toxicsugar22):

that is what i got but i can doit again

OpenStudy (anonymous):

so how did you get a w^6 term there if you added them properly?

OpenStudy (toxicsugar22):

2w^2-4w 2w^2+6w. 2w^2 +2w-1=6w^6+4w-12

OpenStudy (anonymous):

okay, you added 2w^2+2w^2+2w^2 and got 6w^6?

OpenStudy (toxicsugar22):

can u help me please

OpenStudy (anonymous):

you added -4w+6w+2w and got 4w correct?

OpenStudy (anonymous):

then how did you add 2w^2+2w^2+2w^2 and get 6w^6? according to your logic, -4w+6w+2w should be 4w^3 isn't it?

OpenStudy (toxicsugar22):

ohh yes so can u write the answer again please

OpenStudy (anonymous):

no, I can't. I don't think you understood what I am trying to say to you. if you add 2 oranges to 2 oranges, how many oranges do you get?